OpenStudy (anonymous):

Solve: 3(x + 4) = 5(x − 2)

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

First I would distribute the 3 to x+4 and 5 to x - 2

OpenStudy (anonymous):

is this easy for you?

OpenStudy (anonymous):

@camilasf your supposed to walk the person thru the problem with code of conduct.

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

You have to use distributive property to solve this question.

OpenStudy (theviper):

3x+12=5x-10 12+10=5x-3x 22=2x x=22/2=11

OpenStudy (anonymous):

okay thanks!

OpenStudy (theviper):

yw:)

OpenStudy (anonymous):

exactly what i said ... :)

OpenStudy (anonymous):

You will get 3x + 12 = 5x - 10 Subtract 12 from both sides and get 3x = 5x - 22 Then subtract 5x from both sides and get -2x = - 22 and divide by -2 to get x = 11

OpenStudy (anonymous):

okay thanks @TheMusician! Can you come help me with my other?

OpenStudy (anonymous):

What is the problem?

OpenStudy (anonymous):

Solve: -1/2 x + 1 = −x + 8

OpenStudy (anonymous):

I would take the 1 and subtract from both sides to get -1/2 x = −x + 7 then add x to both sides to get 1/2 x = 7, then take 1/2 from both sides to get x = 3 and 1/2 or 7/2

OpenStudy (anonymous):

okay thankss so much!

OpenStudy (anonymous):

No problem D:

OpenStudy (anonymous):

why sad face? D: